# AI-SearchPlanner: Modular Agentic Search via Pareto-Optimal Multi-Objective Reinforcement Learning

### v1

@@ 1 Introduction
Recent research has investigated combining Large Language Models (LLMs) with search engines to harness both the LLMs’ internal knowledge and external information retrieval, thereby enhancing performance on complex reasoning tasks. The early representative paradigm, Retrieval-Augmented Generation (RAG), retrieves relevant passages based on input queries and incorporates them into the LLM’s context for generation. While this approach facilitates the utilization of external knowledge, most RAG systems rely on static retrieval mechanisms, lacking adaptive planning capabilities for complex reasoning. Consequently, their performance remains highly sensitive to retrieval accuracy. With advancements in the general capabilities of LLMs, recent research has explored leveraging search engines as tools — primarily through prompt-based methods or self-supervised fine-tuning (SFT). These approaches enable LLMs to autonomously determine whether and how to invoke search engines based on the problem-solving context. However, prompt-based methods often exhibit limited generalization, particularly for complex reasoning tasks that fall outside the LLM’s pretraining distribution. Although SFT-based methods offer greater adaptability, they require large-scale, high-quality annotated trajectories of search-reasoning interactions, posing scalability challenges. Furthermore, recent studies indicate that SFT methods tend to overfit to memorized reasoning paths, further constraining their generalization to novel or out-of-distribution scenarios.
@@
    Pareto Optimization of Planning Utility and Cost: In real-world scenarios, the trade-off between search planning effectiveness (e.g., end-to-end QA accuracy) and computational overhead (e.g., search frequency, reasoning turns) significantly impacts user experience. We formalize search planning as a pareto optimization problem to jointly maximize planning utility while minimizing planning cost.
@@ 2 Related Work
With the advancement of the general capabilities of LLMs, recent research Qu et al. (2025); Trivedi et al. (2022a); Yao et al. (2023); Schick et al. (2023) has explored integrating search engines as tools to enhance problem-solving. Specially, prompt-based methods Trivedi et al. (2022a); Yao et al. (2023) and self-supervised fine-tuning (SFT) approaches Schick et al. (2023) have been proposed to enable LLMs to autonomously determine when and how to invoke search engines during reasoning. For instance, IRCoT Trivedi et al. (2022a) and ReAct Yao et al. (2023) utilize iterative prompting strategies to interleave reasoning with search engine calls, while Toolformer Schick et al. (2023) leverages supervised fine-tuning to improve search-related capabilities. However, prompt-based methods often exhibit limited generalization, especially for complex reasoning tasks that lie outside the pretraining distribution of LLMs. SFT-based methods often depend on large-scale, high-quality annotated trajectories of search-reasoning interactions, which limits their scalability. Moreover, recent studies suggest that SFT-based approaches tend to overfit to memorized reasoning paths, further constraining their generalization to novel or out-of-distribution scenarios. Recent studies Jin et al. (2025); Song et al. (2025); Wu et al. (2025a); Li et al. (2025a) demonstrate that reinforcement learning (RL) can enable LLMs to develop advanced reasoning capabilities using only outcome-based rewards. For example, Search-R1-like approaches Jin et al. (2025); Song et al. (2025) have investigated the potential of RL methods such as PPO Schulman et al. (2017) and GRPO Shao et al. (2024) in agentic search scenarios. By employing outcome-based reward mechanisms and end-to-end training, these methods enable LLMs to jointly optimize multi-turn search planning and answer generation, achieving strong performance in QA tasks. WebDancer Wu et al. (2025a) and WebSailor Li et al. (2025a) enhance LLMs for autonomous web tasks through data-centric training but employ distinct methodologies. WebDancer follows a four-stage paradigm: (1) curating browsing data, (2) sampling trajectories, (3) supervised fine-tuning, and (4) reinforcement learning, leveraging ReAct for reasoning-action interleaving. In contrast, WebSailor generates high-uncertainty tasks via structured sampling and obfuscation, initializing with RFT and training with DUPO.
@@ 3 AI-SearchPlanner
In this section, we introduce AI-SearchPlanner, a novel reinforcement learning framework designed to enhance end-to-end QA performance by focusing on search planning. As illustrated in Figure 1, we demonstrate AI-SearchPlanner framework. Given a question $q$ and its ground truth answer $gt$, a trainable search planner $LLM_{plan}$ iteratively interacts with a search engine $S\left(\cdot\right)$: At each turn $t$, $LLM_{plan}$ first generates reasoning context, and then decides to either: 1) Generates sub-queries $\left\{sq\right\}^{t}$ within search tool to retrieve relevant documents via $S\left(\cdot\right)$, or 2) Terminates reasoning by generating call_answer_llm tool , which feeds the accumulated knowledge from the reasoning trajectory into a frozen generator $LLM_{gen}$ to produce an answer $a$. The final multi-turn reasoning trajectory is represented as $T$ = $\left\{q,r^{S}_{1},S\left(\left\{sq\right\}^{1}\right),...,r^{S}_{L},S\left(\left\{sq\right\}^{L}\right),r^{A}\right\}$, where $r^{S}_{\ast}$ and $r^{A}$ refer to the planning contexts of searching sub-queries and invoking a frozen generator, respectively.
In the following, we introduces three key innovations in AI-SearchPlanner to achieve the objectives: 1) Decoupling the Architecture of the Search Planner and Generator, 2) Dual-Reward Alignment for Search Planning, and 3) Pareto Optimization of Planning Utility and Cost.
@@ 3.4 Reinforcement Learning Training
  • Optimization of Search Planner. We optimize the search planning policy using reinforcement learning with a dual reward formulation. Each rollout consists of a complete search trajectory, encompassing the input question, a sequence of retrieved documents, and a termination decision (i.e., call generator to answer). Upon constructing the final context, a frozen generator generates the answer, and the dual reward is computed. In this work, we employ Proximal Policy Optimization (PPO) Schulman et al. (2017). The corresponding PPO objective is defined as:

### v3

@@ 1 Introduction
Recent research has investigated combining Large Language Models (LLMs) with search engines to harness both the LLMs’ internal knowledge and external information retrieval (IR), thereby enhancing performance on complex reasoning tasks. The early representative paradigm, Retrieval-Augmented Generation (RAG), retrieves relevant passages based on input queries and incorporates them into the LLM’s context for generation. While this approach facilitates the utilization of external knowledge, most RAG systems rely on static retrieval mechanisms, lacking adaptive planning capabilities for complex reasoning. Consequently, their performance remains highly sensitive to retrieval accuracy. With advancements in the general capabilities of LLMs, recent research has explored leveraging search engines as tools — primarily through prompt-based methods or self-supervised fine-tuning (SFT). These approaches enable LLMs to autonomously determine whether and how to invoke search engines based on the problem-solving context. However, prompt-based methods often exhibit limited generalization, particularly for complex reasoning tasks that fall outside the LLM’s pretraining distribution. Although SFT-based methods offer greater adaptability, they require large-scale, high-quality annotated trajectories of search-reasoning interactions, posing scalability challenges. Furthermore, recent studies indicate that SFT methods tend to overfit to memorized reasoning paths, further constraining their generalization to out-of-distribution scenarios.
@@
    Pareto Optimization of Planning Utility and Cost: In real-world scenarios, the trade-off between search planning effectiveness (e.g., end-to-end QA accuracy) and computational overhead (e.g., reasoning turns) significantly impacts user experience. We formalize search planning as a pareto optimization problem to jointly maximize planning utility while minimizing planning cost.
@@ 2 Related Work
With the advancement of the general capabilities of LLMs, recent research Qu et al. (2025); Trivedi et al. (2022a); Yao et al. (2023); Schick et al. (2023) has explored integrating search engines as tools to enhance problem-solving. Specially, prompt-based methods Trivedi et al. (2022a); Yao et al. (2023) and self-supervised fine-tuning (SFT) approaches Schick et al. (2023) have been proposed to enable LLMs to autonomously determine when and how to invoke search engines during reasoning. For instance, IRCoT Trivedi et al. (2022a) and ReAct Yao et al. (2023) utilize iterative prompting strategies to interleave reasoning with search engine calls, while Toolformer Schick et al. (2023) leverages supervised fine-tuning to improve search-related capabilities. However, prompt-based methods often exhibit limited generalization, especially for complex reasoning tasks that lie outside the pretraining distribution of LLMs. SFT-based methods often depend on large-scale, high-quality annotated trajectories of search-reasoning interactions, which limits their scalability. Moreover, recent studies suggest that SFT-based approaches tend to overfit to memorized reasoning paths, further constraining their generalization to novel or out-of-distribution scenarios. Recent studies Jin et al. (2025); Song et al. (2025); Wu et al. (2025a); Li et al. (2025a; c) demonstrate that reinforcement learning (RL) can enable LLMs to develop advanced reasoning capabilities using only outcome-based rewards. For example, Search-R1-like approaches Jin et al. (2025); Song et al. (2025) have investigated the potential of RL methods such as PPO Schulman et al. (2017) and GRPO Shao et al. (2024) in agentic search scenarios. By employing outcome-based reward mechanisms and end-to-end training, these methods enable LLMs to jointly optimize multi-turn search planning and answer generation, achieving strong QA performance. A related concurrent work is s3 Jiang et al. (2025), which also decouples the planner from the generator. Our approach differs from s3 as follows: 1) Objective: s3 aims to train a high-performing planner with minimal data, while we focuses on industrial applications, optimizing the planner for the best utility-cost trade-off using state-of-the-art LLMs. 2) Framework: s3 provides the generator with key documents from the final reasoning step; we instead supply the entire planning trajectory. 3) Method Design: s3 introduces a Gain Beyond RAG reward and filters out training examples solvable by naive RAG, thereby improving end-to-end QA performance on complex questions. We directly optimizes planning ability with an outcome-process dual reward and incorporates Pareto optimization of utility and cost, significantly reducing cost while maintaining performance. Additionally, another similar line of recent work Xiao et al. (2024); Shao et al. (2025), ReasoningIR, enhance IR models with intrinsic reasoning ability, by using LLM reasoning to generate query-document pairs for training. In contrast, we leverages query–search–document planning trajectories to directly produce high-quality answers.
@@ 3 AI-SearchPlanner
In this section, we introduce AI-SearchPlanner, a novel reinforcement learning framework designed to enhance end-to-end QA performance by focusing on search planning. As illustrated in Figure 1, we demonstrate AI-SearchPlanner framework. Given a question $q$ and its ground truth answer $gt$, a trainable search planner $LLM_{plan}$ iteratively interacts with a search engine $S\left(\cdot\right)$: At each turn $t$, $LLM_{plan}$ first generates reasoning context, and then decides to either: 1) Generates sub-queries $\left\{sq\right\}^{t}$ within search tool to retrieve relevant documents via $S\left(\cdot\right)$, or 2) Terminates reasoning by generating call_answer_llm tool , which feeds the accumulated knowledge from the reasoning trajectory into a frozen generator $LLM_{gen}$ to produce an answer $a$. The final multi-turn reasoning trajectory is represented as $T$ = $\left\{q,r^{S}_{1},S\left(\left\{sq\right\}^{1}\right),...,r^{S}_{L},S\left(\left\{sq\right\}^{L}\right),r^{A}\right\}$, where $r^{S}_{\ast}$ and $r^{A}$ refer to the planning contexts of searching sub-queries and invoking a frozen generator, respectively. In the following, we introduces three key innovations in AI-SearchPlanner to achieve the objectives.
@@ 3.4 Reinforcement Learning Training
  • Optimization of Search Planner. We optimize the search planning policy using reinforcement learning with a dual reward formulation. Each search planning trajectory comprises an input question, a sequence of retrieved documents, and a termination decision (i.e., call generator to answer). Upon constructing the final context, a frozen generator generates the answer, and the dual reward is computed. In this work, we employ Proximal Policy Optimization (PPO) Schulman et al. (2017). The corresponding PPO objective is defined as: